# Graph Positional Autoencoders as Self-supervised Learners

### v1

@@ Abstract.
Graph self-supervised learning seeks to learn effective graph representations without relying on labeled data. Among various approaches, graph autoencoders (GAEs) have gained significant attention for their efficiency and scalability. Typically, GAEs take incomplete graphs as input and predict missing elements, such as masked nodes or edges. While effective, our experimental investigation reveals that traditional node or edge masking paradigms primarily capture low-frequency signals in the graph and fail to learn the expressive structural information. To address these issues, we propose Graph Positional Autoencoders (GraphPAE), which employs a dual-path architecture to reconstruct both node features and positions. Specifically, the feature path uses positional encoding to enhance the message-passing processing, improving GAE’s ability to predict the corrupted information. The position path, on the other hand, leverages node representations to refine positions and approximate eigenvectors, thereby enabling the encoder to learn diverse frequency information. We conduct extensive experiments to verify the effectiveness of GraphPAE, including heterophilic node classification, graph property prediction, and transfer learning. The results demonstrate that GraphPAE achieves state-of-the-art performance and consistently outperforms baselines by a large margin.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3736990ccs: Information systems Data mining
@@
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15516721.
@@ 1. Introduction
(c) Eigenvector offset
Figure 1. Frequency magnitude comparison between the original graph and its corrupted versions.
@@
Among existing approaches, the generative (Wu et al., 2021b; Waikhom and Patgiri, 2023) and contrastive (Ju et al., 2024) learning paradigms have dominated recent advances. In particular, graph autoencoders (GAEs) (Kipf and Welling, 2016; Hou et al., 2022) have regained attention due to their simplicity, efficiency, and scalability. The GAEs follow a corruption-reconstruction framework, which learns graph representations by recovering the missing information of the incomplete input graphs, as Table 1 illustrates. For example, GraphMAE (Hou et al., 2022) replaces node features with a learnable token, while Bandana (Zhao et al., 2024) proposes a non-discrete edge masking strategy. Moreover, some GAEs even go beyond reconstructing nodes and edges by targeting structural features like degree (Li et al., 2023a). Despite their success, the performance of GAEs highly depends on the choice of corruption and reconstruction objectives. Therefore, a natural question arises: Do existing feature or edge masking mechanisms enable GAEs to sufficiently capture key structural patterns in graph data? If not, what alternative objectives could be designed to further improve the performance of GAEs? A well-informed answer can not only help us identify the weaknesses of existing masking strategies but also deepen our understanding of GAEs.
@@
To answer this question, we first revisit the feature and edge masking strategies from a spectral perspective (Bo et al., 2023b, a). Specifically, we transform the node features $\mathbf{X}$ into the spectral domain by using the eigenvectors of graph Laplacian $\mathbf{U}^{\top}$ and examine its frequency magnitude $\mathbf{U}^{\top}\mathbf{X}$. In the case of feature masking, we randomly mask 20% node features in the Squirrel dataset, denoting as $\widetilde{\textbf{X}}$. The comparison between $\mathbf{U}^{\top}\mathbf{X}$ and $\mathbf{U}^{\top}\widetilde{\textbf{X}}$ is shown in Figure 1(a). For edge masking, we randomly remove 20% edges and construct the corrupted graph eigenvectors $\widetilde{\textbf{U}}^{\top}$. Figure 1(b) compares $\mathbf{U}^{\top}\mathbf{X}$ and $\widetilde{\textbf{U}}^{\top}\mathbf{X}$. Experimental details are provided in Appendix A.1. From these comparisons, we observe that the magnitude differences between original and corrupted graphs are more pronounced in the low-frequency band [0.0-0.1], while the differences decrease at higher frequencies. This indicates that existing GAEs focus primarily on reconstructing low-frequency information. Specifically, the significant difference in the [0.0-0.1] band suggests that masking predominantly damage these low-frequency components. During reconstruction, the GAE is optimized by minimizing the reconstruction loss, which inherently penalizes larger discrepancies more heavily to accelerate convergence. Since low-frequency components are corrupted more severely, GAE implicitly tends to prioritize these components to reduce the loss quickly. As a result, existing GAEs often overlook high-frequency information, which has been shown to be valuable in real-world tasks, such as heterophilic node classification (Bo et al., 2021; Wang et al., 2020).
@@
Once the weakness of existing GAEs is identified, it is natural to ask: How can we design the corruption and reconstruction objectives to enable GAEs to learn diverse frequency information? Essentially, the eigenvectors of the graph Laplacian represent different frequencies. Motivated by this, a straightforward approach is to perturb eigenvectors. Figure 1(c) illustrates the impact on frequency magnitudes of corrupted eigenvectors $\widetilde{\mathbf{U}}$ with offsets. By comparing $\mathbf{U}^{\top}\mathbf{X}$ and $\widetilde{\mathbf{U}}^{\top}\mathbf{X}$, we observe that the eigenvector offset corrupts higher-frequency signals and covers a broader range. These observations indicate that the eigenvector offset introduces wider perturbations than masking, enabling the model to capture more diverse information. However, this task is non-trivial and presents two main challenges: (1) Eigenvectors represent the underlying structures of a graph, which cannot be easily approximated by basic GNNs. Existing GAEs typically adopt message-passing neural networks (MPNNs) as the encoder, whose expressiveness is bounded by the 1-WL test (Xu et al., 2018). The intrinsic weakness of MPNNs restricts GAEs’ ability to capture long-range dependencies between nodes and higher-order graph patterns (Wu et al., 2021a; Dwivedi et al., 2022). (2) Eigenvectors suffer from sign- and basis-ambiguity issues (Lim et al., 2022). Directly reconstructing the eigenvectors leads to non-unique solutions, affecting the robustness of GAEs (Wang et al., 2022; Bo et al., 2024).
@@
To overcome these challenges, we propose Graph Positional Autoencoders (GraphPAE) for graph self-supervised learning. GraphPAE adopts a dual-path architecture to address both the expressivity and ambiguity issues. Specifically, in the feature path, GraphPAE integrates positional encoding (PE) into the message-passing process to enhance the expressiveness of MPNNs (Dwivedi et al., 2023; Wang et al., 2022), thus improving the model’s ability to reconstruct corrupted information. Moreover, in the position path, node representations are used to refine the PE, which approximates the eigenvectors, thereby transferring diverse frequency information into the encoder. Finally, in the reconstruction stage, instead of directly recovering the raw eigenvectors, GraphPAE uses the relative node distance as a surrogate objective, avoiding potential ambiguity issues.
@@
    To the best of our knowledge, we are the first to explore the masking strategy in GAEs from a spectral perspective. By comparing the frequency magnitudes between original and corrupted graphs, we identify that existing GAEs focus on reconstructing the extremely low-frequency information of graphs and neglecting other frequencies.
@@
    We propose GraphPAE, a novel graph positional autoencoder that reconstructs both node features and positions, thereby enabling GAEs to capture a broader range of frequency information. GraphPAE adopts a dual-path architecture. Specifically, the feature path generates expressive node presentations enhanced by positional encodings, and the position path outputs positional encodings to facilitate position reconstruction.
@@ 2.1. Generative Graph Self-supervised Learning
Graph Autoregressive models (GARs) treat sequential graph generation as the pre-training task, where the node or edge is predicted based on its prior context. GPT-GNN (Hu et al., 2020a) factorizes each node generation into attribute and edge generation and replenishes the omitted parts by an adaptive queue. MGSSL (Zhang et al., 2021a) introduces motif generation into pre-training. GraphsGPT (Gao et al., [n. d.]) transforms non-Euclidean graphs into learnable Euclidean words and reconstructs the original graph from these words.
Graph Autoencoders (GAEs) reconstruct all desired content once from the latent representation output by the encoder. Early GAEs (e.g., VGAE (Kipf and Welling, 2016), ARGVA (Pan et al., 2018)) learn representations through link reconstruction and spark a series of work, including feature reconstruction (e.g., GALA (Park et al., 2019), WGDN (Cheng et al., 2023)) and combination reconstruction of structure and feature (e.g., GATE (Salehi and Davulcu, 2019)). However, these traditional GAEs often perform poorly in downstream tasks, except link prediction, which is attributed to their overemphasis on proximity information at the expense of structural information (Hassani and Khasahmadi, 2020). Recently, masked GAEs (Hu et al., [n. d.]; Hou et al., 2022) have become highly successful models for representation learning. GraphMAE (Hou et al., 2022) and GraphMAE2 (Hou et al., 2023) successfully bridge the performance gap between graph contrastive learning and generative learning by reconstructing masked features for training. AUG-MAE (Wang et al., 2024) introduces an adversarial masking strategy to enhance feature alignment and add a uniformity regularizer to promote high-quality graph representations. Additionally, some works are focusing on masking and reconstructing graph structures (e.g., edges (Li et al., 2023a; Tan et al., 2023) and paths (Li et al., 2023a)). Bandana (Zhao et al., 2024) uses continuous and dispersive edge masks and bandwidth prediction instead of discrete edge masks and reconstruction. Besides, some works mask features and edges simultaneously (Li et al., 2023b; Shi et al., 2023) and propose novel reconstruction objectives such as latent embeddings (Shi et al., 2023).
@@ 2.2. Graph Positional Encoding
Laplacian-based PEs leverage the eigenvectors of the graph Laplacian matrix as initial node positional features (Dwivedi and Bresson, 2020; Dwivedi et al., 2023, 2021). Eigenvectors form a natural basis for encoding graph structure, similar to how sinusoidal functions represent positional information in sequential models (Vaswani, 2017). Serving as PEs, eigenvectors have two traditional constraints, i.e., sign- and basis ambiguity (Lim et al., 2022). To keep the models robust to sign ambiguity, (Dwivedi and Bresson, 2020; Dwivedi et al., 2023) randomly flip the sign of the eigenvectors during training. SAN (Kreuzer et al., 2021) introduces the information of eigenvalues into PEs and leverages a transformer-based positional encoder, enabling models to be more informative and expressive. PEG (Wang et al., 2022) solves the sign- and basis-ambiguity by treating the distance of eigenvectors between node pairs as PEs. (Lim et al., 2022) propose SignNet and BasisNet to learn sign- and basis-invariant PEs, respectively. (Huang et al., 2023; Bo et al., 2024) study stable PEs that are robust to disturbance of the Laplacian matrix.
Distance-based PEs assign positional features to nodes based on their relative distances within the graph structure, which are typically derived from spatial relationships (Rampášek et al., 2022; Li et al., 2020). One common approach is to use the random walk matrix capturing structural relationships in a probabilistic manner (Dwivedi et al., 2021, 2023). Another popular method is to encode the shortest distance between node pairs (Ying et al., 2021; Rampášek et al., 2022; Li et al., 2020). For example, Graphormer (Ying et al., 2021) incorporates shortest-path distances into its attention mechanism. Additionally, GraphiT (Mialon et al., 2021) encodes PEs with a diffusion kernel, enabling more flexible and adaptive representations of node relationships. (Shiv and Quirk, 2019) introduces a novel positional encoding scheme that extends transformers to tree-structured data, enabling efficient and parallelizable sequence-to-tree, tree-to-sequence, and tree-to-tree mappings. (Wijesinghe et al., [n. d.]) first introduces PEs in graph self-supervised learning and designs a GNN framework that integrates a k-hop message-passing mechanism to enhance its expressiveness.
@@
Figure 2. (a): GraphPAE integrates a positional corruption-reconstruction mechanism to encourage the GAE to capture diverse frequency information. For feature reconstruction, masked features are encoded with $\mathbf{P}$ and decoded to recover original features. For positional reconstruction, noise is added to $\widetilde{\mathbf{U}}$ to corrupt relative distances, which are then encoded with $\mathbf{X}$ and decoded to recover the original distances. Figure (b): The encoder employs a dual-path architecture to update node $\mathbf{X}^{\left(l\right)}$ and positional representations $\mathbf{P}^{\left(l\right)}$ at each layer. The feature path integrates positional encodings to enhance message passing, improving GAE’s ability to reconstruct corrupted features. The position path utilizes node representations to refine positional embeddings to approximate original pairwise distances.
@@ 3. Preliminaries
Problem Definition. Given a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$, where $\mathcal{V}$ is the node set with $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$, $\mathcal{E}$ is the set of edges, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ represents the $d$-dimensional node feature matrix. We use the adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$ to describe the graph structure, where $A_{ij}=1$ if there is en edge between $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise. The goal of our framework is to learn an effective graph encoder $f_{enc}(\cdot)$ without relying on labels from downstream tasks. Once trained, the encoder $f_{enc}(\cdot)$ generates node representations as $\mathbf{H}=f_{enc}(\mathcal{G})\in\mathbb{R}^{N\times d_{h}}$ or graph representations as $\mathbf{H}_{\mathcal{G}}=\text{READOUT}\left(\mathbf{H}_{v_{i}}\mid v_{i}\in\mathcal{V}\right)\in\mathbb{R}^{d_{h}}$, where READOUT is a permutation-invariant function such as mean, max, or sum pooling. The representations can then be used to train a simple linear classifier on labeled data from downstream tasks.
Laplacian Graph Eigenvectors. The normalized graph Laplacian matrix $\mathbf{L}$ is defined as $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$, where $\mathbf{I}_{N}$ is an identity matrix and $\mathbf{D}$ is the degree matrix with $D_{ii}=\sum_{j}A_{ij}$ for $v_{i}\in\mathcal{V}$, and $D_{ij}=0$ for $i\neq j$. The Laplacian matrix $\mathbf{L}$ can be decomposed as $\mathbf{L}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$, where $\bm{\Lambda}=\text{diag}(\{\lambda_{i}\}_{i=1}^{N})$ is the diagonal matrix of eigenvalues, and $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{N}]\in\mathbb{R}^{N\times N}$ consists of a set of eigenvectors. Each eigenvector $\mathbf{u}_{i}\in\mathbb{R}^{N}$ corresponds to eigenvalue $\lambda_{i}$, where the eigenvalues are ordered as $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$. Given eigen-decomposition $\mathbf{L}\mathbf{u}_{i}=\lambda_{i}\mathbf{u}_{i}$ with $\mathbf{u}_{i}^{\top}\mathbf{u}_{i}=1$, we have $\mathbf{u}_{i}^{\top}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$. Since $\mathbf{u}_{i}^{\top}\mathbf{L}\mathbf{u}_{i}=\sum_{(j,k)\in\mathcal{E}}(\mathbf{u}_{i,j}-\mathbf{u}_{i,k})^{2}$, we finally derive that $\sum_{(j,k)\in\mathcal{E}}(\mathbf{u}_{i,j}-\mathbf{u}_{i,k})^{2}=\lambda_{i}$, which indicates that $\lambda_{i}$ reflects the frequency magnitude of eigenvector $\mathbf{u}_{i}$ over the graph.
In this paper, we follow the popular positional encoding methods and adopt top-$k$ eigenvectors as initial positions to reduce complexity. Therefore, we redefine $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{K}]\in\mathbb{R}^{N\times K}$, where $K\leq N$. Notably, these eigenvectors are not unique, as sign flips or coordinate rotations may still yield valid decompositions (Wang et al., 2022; Lim et al., 2022). Such sign and basis ambiguity may cause inconsistency in PEs, leading to unstable representation learning in GNNs. Therefore, models should be invariant to these ambiguities for robust performance.

### v2

@@ Abstract.
Graph self-supervised learning seeks to learn effective graph representations without relying on labeled data. Among various approaches, graph autoencoders (GAEs) have gained significant attention for their efficiency and scalability. Typically, GAEs take incomplete graphs as input and predict missing elements, such as masked node features or edges. Although effective, our experimental investigation reveals that traditional feature or edge masking paradigms primarily capture low-frequency signals in the graph and fail to learn expressive structural information. To address these issues, we propose Graph Positional Autoencoders (GraphPAE), which employ a dual-path architecture to reconstruct both node features and positions. Specifically, the feature path uses positional encoding to enhance the message-passing processing, improving the GAEs’ ability to predict the corrupted information. The position path, on the other hand, leverages node representations to refine positions and approximate eigenvectors, thereby enabling the encoder to learn diverse frequency information. We conduct extensive experiments to verify the effectiveness of GraphPAE, including heterophilic node classification, graph property prediction, and transfer learning. The results demonstrate that GraphPAE achieves state-of-the-art performance and consistently outperforms the baselines by a large margin.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736990isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Data mining
@@
The source code of this paper has been made publicly available at %****␣sample-sigconf.tex␣Line␣275␣****https://doi.org/10.5281/zenodo.15516721.
@@ 1. Introduction
(c) Eigenvector offsetting
Figure 1. Frequency magnitudes of the original and corrupted graphs in the Squirrel dataset.
@@
Among existing approaches, the generative (Wu et al., 2021b; Waikhom and Patgiri, 2023) and contrastive (Ju et al., 2024) learning paradigms have dominated recent advances. In particular, graph autoencoders (GAEs) (Kipf and Welling, 2016; Hou et al., 2022) have gained attention due to their simplicity, efficiency, and scalability. The GAEs follow a corruption-reconstruction framework, which learns graph representations by recovering the missing information of the incomplete input graphs, as Table 1 illustrates. For example, GraphMAE (Hou et al., 2022) replaces node features with a learnable token, while Bandana (Zhao et al., 2024) proposes a non-discrete edge masking strategy. Moreover, some GAEs even go beyond reconstructing node features and edges by targeting structural features like degree (Li et al., 2023a). Despite their success, the performance of GAEs highly depends on the choice of corruption and reconstruction objectives. Therefore, a natural question arises: Do existing feature or edge masking mechanisms fully exploit the graph data? If not, what alternative objectives could be designed to further improve the performance of GAEs? A well-informed answer can help us identify the weaknesses of existing masking strategies and deepen our understanding of GAEs.
@@
To answer this question, we first revisit the feature and edge masking strategies from a spectral perspective (Bo et al., 2023b, a). Specifically, we transform the node features $\mathbf{X}$ into the spectral domain by using the eigenvectors of graph Laplacian $\mathbf{U}^{\top}$ and examine its frequency magnitude $\mathbf{U}^{\top}\mathbf{X}$. In case of feature masking, we randomly mask 20% node features and denote it as $\widetilde{\textbf{X}}$, while for edge masking, we randomly remove 20% edges and construct the corrupted graph eigenvectors $\widetilde{\textbf{U}}^{\top}$. The frequency magnitude changes of $\mathbf{U}^{\top}\widetilde{\textbf{X}}$ and $\widetilde{\textbf{U}}^{\top}\mathbf{X}$ are shown in Figures 1(a) and 1(b), respectively. It can be observe that the magnitude differences between original and corrupted graphs are more pronounced in the low-frequency band, i.e., [0.0, 0.1], while the differences decrease at higher frequencies. Notably, to effectively minimize the reconstruction loss, GAEs will primarily focus on the frequency bands with larger discrepancies. As a result, existing GAEs mainly reconstruct low-frequency information and overlook the high-frequency information, which has been shown to be valuable in real-world tasks (Bo et al., 2021; Wang et al., 2020). More experimental details are provided in Appendix A.
@@
Given the above weakness of existing GAEs, it is natural to ask: How can we design the corruption and reconstruction objectives to exploit the diverse frequency information? Essentially, the eigenvectors of graph Laplacian represent different frequencies. Directly perturbing and reconstructing the eigenvectors can help GAEs learn different frequency information in graphs. Figure 1(c) illustrates the frequency magnitudes of the eigenvectors with random offsets, i.e., $\widetilde{\mathbf{U}}^{\top}X=(\mathbf{U}+\delta)^{\top}X$. We observe that this strategy perturbs a broader range of frequency components, particularly in higher-frequency, i.e. [0.2, 0.5], enabling the model to capture more diverse frequency information. However, reconstructing the eigenvectors is not a trivial task and presents two main challenges: (1) Eigenvectors represent the global structural patterns of a graph, which cannot be easily approximated by basic GNNs. Existing GAEs typically adopt message-passing neural networks (MPNNs) as the encoder, whose expressiveness is bounded by the 1-WL test (Xu et al., 2018). The intrinsic weakness of MPNNs restricts GAEs’ ability to capture long-range dependencies between nodes and higher-order graph patterns (Wu et al., 2021a; Dwivedi et al., 2022). (2) Eigenvectors suffer from sign- and basis-ambiguity issues (Lim et al., 2022). Directly reconstructing the eigenvectors leads to non-unique solutions, affecting the robustness of GAEs (Wang et al., 2022; Bo et al., 2024).
@@
To overcome these challenges, we propose Graph Positional Autoencoders (GraphPAE) for graph self-supervised learning. GraphPAE adopts a dual-path architecture to address both the expressivity and ambiguity issues. Specifically, in the feature path, GraphPAE integrates positional encoding (PE) into the message-passing process to enhance the expressiveness of MPNNs, thus improving the model’s ability to reconstruct corrupted information. Moreover, in the position path, node representations are used to refine the PE, which approximates the eigenvectors, thereby transferring diverse frequency information into the encoder. Finally, in the reconstruction stage, instead of directly recovering the raw eigenvectors, GraphPAE uses the relative node distance as a surrogate objective, avoiding potential ambiguity issues.
@@
    We are the first to explore the masking strategy in GAEs from a spectral perspective. By comparing the frequency magnitudes between original and corrupted graphs, we identify that existing GAEs focus on reconstructing the low-frequency information of graphs and neglecting other frequencies.
@@
    We propose GraphPAE, a novel graph positional autoencoder that learns graph representations by reconstructing both node features and positions, thereby enabling GAEs to capture a broader range of frequency information.
@@ 2. Related Work
We provide an overview of two main areas related to our work, namely generative approaches in graph self-supervised learning and positional encoding techniques to enhance structural expressiveness in GNNs.
@@ 2.1. Generative Graph Self-supervised Learning
Graph Autoregressive models (GARs). GARs treat sequential graph generation as the pre-training task, where the node or edge is predicted based on its prior context. GPT-GNN (Hu et al., 2020a) factorizes each node generation into attribute and edge generation and replenishes the omitted parts by an adaptive queue. MGSSL (Zhang et al., 2021a) introduces motif generation into pre-training. GraphsGPT (Gao et al., [n. d.]) transforms non-Euclidean graphs into learnable Euclidean words and reconstructs the original graph from these words.
Graph Autoencoders (GAEs). GAEs reconstruct all desired content once from the latent representation output by the encoder. Early GAEs (e.g., VGAE (Kipf and Welling, 2016) and ARGVA (Pan et al., 2018)) learn representations through link reconstruction and spark a series of work, including feature reconstruction (e.g., GALA (Park et al., 2019) and WGDN (Cheng et al., 2023)) and combination reconstruction of structure and feature (e.g., GATE (Salehi and Davulcu, 2019)). However, these traditional GAEs often perform poorly in downstream tasks, except link prediction, which is attributed to their overemphasis on proximity information at the expense of structural information (Hassani and Khasahmadi, 2020). Recently, masked GAEs (Hu et al., [n. d.]; Hou et al., 2022) have become highly successful models for representation learning. GraphMAE (Hou et al., 2022) and GraphMAE2 (Hou et al., 2023) successfully bridge the performance gap between graph contrastive learning and generative learning by reconstructing masked features for training. AUG-MAE (Wang et al., 2024) introduces an adversarial masking strategy to enhance feature alignment and add a uniformity regularizer to promote high-quality graph representations. Additionally, some works are focusing on masking and reconstructing graph structures (e.g., edges (Li et al., 2023a; Tan et al., 2023) and paths (Li et al., 2023a)). Bandana (Zhao et al., 2024) uses continuous and dispersive edge masks and bandwidth prediction instead of discrete edge masks and reconstruction. Besides, some works mask features and edges simultaneously (Li et al., 2023b; Shi et al., 2023) and propose novel reconstruction objectives such as latent embeddings (Shi et al., 2023).
@@ 2.2. Graph Positional Encoding
Laplacian-based PEs. These methods leverage the eigenvectors of the graph Laplacian matrix as initial node positional features (Dwivedi and Bresson, 2020; Dwivedi et al., 2023, 2021). Eigenvectors form a natural basis for encoding graph structure, similar to how sinusoidal functions represent positional information in sequential models (Vaswani, 2017). Serving as PEs, eigenvectors have two traditional constraints, i.e., sign- and basis ambiguity (Lim et al., 2022). To keep the models robust to sign ambiguity, (Dwivedi and Bresson, 2020; Dwivedi et al., 2023) randomly flip the sign of the eigenvectors during training. SAN (Kreuzer et al., 2021) introduces the information of eigenvalues into PEs and leverages a transformer-based positional encoder, enabling models to be more informative and expressive. PEG (Wang et al., 2022) solves the sign- and basis-ambiguity by treating the distance of eigenvectors between node pairs as PEs. (Lim et al., 2022) propose SignNet and BasisNet to learn sign- and basis-invariant PEs, respectively. (Huang et al., 2023; Bo et al., 2024) study stable PEs that are robust to disturbance of the Laplacian matrix.
Distance-based PEs. These approaches assign positional features to nodes based on their relative distances within the graph structure, which are typically derived from spatial relationships (Rampášek et al., 2022; Li et al., 2020). One common approach is to use the random walk matrix capturing structural relationships in a probabilistic manner (Dwivedi et al., 2021, 2023). Another popular method is to encode the shortest distance between node pairs (Ying et al., 2021; Rampášek et al., 2022; Li et al., 2020). For example, Graphormer (Ying et al., 2021) incorporates shortest-path distances into its attention mechanism. Additionally, GraphiT (Mialon et al., 2021) encodes PEs with a diffusion kernel, enabling more flexible and adaptive representations of node relationships. (Shiv and Quirk, 2019) introduces a novel positional encoding scheme that extends transformers to tree-structured data, enabling efficient and parallelizable sequence-to-tree, tree-to-sequence, and tree-to-tree mappings. (Wijesinghe et al., [n. d.]) first introduces PEs in graph self-supervised learning and designs a GNN framework that integrates a k-hop message-passing mechanism to enhance its expressiveness.
@@
Figure 2. (a): GraphPAE integrates a positional corruption-reconstruction mechanism to encourage the GAE to capture diverse frequency information. For feature reconstruction, masked features are encoded with $\mathbf{P}$ and decoded to recover original features. For positional reconstruction, noise is added to $\widetilde{\mathbf{U}}$ to corrupt relative distances, which are then encoded with $\mathbf{X}$ and decoded to recover the original distances. Figure (b): The encoder employs a dual-path architecture to update node $\mathbf{X}^{\left(l\right)}$ and positional representations $\mathbf{P}^{\left(l\right)}$ at each layer. The feature path integrates positional encodings to enhance message passing, improving the GAE’s ability to reconstruct corrupted features. The position path utilizes node representations to refine positional embeddings to approximate original pairwise distances.
@@ 3. Preliminaries
Problem Definition. Given a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$, where $\mathcal{V}$ is the node set with $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$, $\mathcal{E}$ is the set of edges, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ represents the $d$-dimensional node feature matrix. We use the adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$ to describe the graph structure, where $A_{ij}=1$ if there is an edge between $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise. The goal of our framework is to learn an effective graph encoder $f_{enc}(\cdot)$ without relying on labels from downstream tasks. Once trained, the encoder $f_{enc}(\cdot)$ generates node representations as $\mathbf{H}=f_{enc}(\mathcal{G})\in\mathbb{R}^{N\times d_{h}}$ or graph representations as $\mathbf{H}_{\mathcal{G}}=\text{READOUT}\left(\mathbf{H}_{v_{i}}\mid v_{i}\in\mathcal{V}\right)\in\mathbb{R}^{d_{h}}$, where READOUT is a permutation-invariant function such as mean, max or sum pooling. The representations can then be used to train a simple linear classifier on labeled data from downstream tasks.
Laplacian Graph Eigenvectors. The normalized graph Laplacian matrix $\mathbf{L}$ is defined as $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$, where $\mathbf{I}_{N}$ is an identity matrix and $\mathbf{D}$ is the degree matrix with $D_{ii}=\sum_{j}A_{ij}$ for $v_{i}\in\mathcal{V}$, and $D_{ij}=0$ for $i\neq j$. The Laplacian matrix $\mathbf{L}$ can be decomposed as $\mathbf{L}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$, where $\bm{\Lambda}=\text{diag}(\{\lambda_{i}\}_{i=1}^{N})$ is the diagonal matrix of eigenvalues, and $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{N}]\in\mathbb{R}^{N\times N}$ consists of a set of eigenvectors. Each eigenvector $\mathbf{u}_{i}\in\mathbb{R}^{N}$ corresponds to eigenvalue $\lambda_{i}$, where the eigenvalues are ordered as $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$. Given eigen-decomposition $\mathbf{L}\mathbf{u}_{i}=\lambda_{i}\mathbf{u}_{i}$ with $\mathbf{u}_{i}^{\top}\mathbf{u}_{i}=1$, we have $\mathbf{u}_{i}^{\top}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$. Since $\mathbf{u}_{i}^{\top}\mathbf{L}\mathbf{u}_{i}=\sum_{(j,k)\in\mathcal{E}}(\mathbf{u}_{i,j}-\mathbf{u}_{i,k})^{2}$, we finally derive that $\sum_{(j,k)\in\mathcal{E}}(\mathbf{u}_{i,j}-\mathbf{u}_{i,k})^{2}=\lambda_{i}$, which indicates that $\lambda_{i}$ reflects the frequency magnitude of eigenvector $\mathbf{u}_{i}$ over the graph. In this paper, we follow the popular positional encoding methods and adopt top-$k$ eigenvectors as initial positions to reduce complexity. Therefore, we redefine $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{K}]\in\mathbb{R}^{N\times K}$, where $K\leq N$.